Three friends started a part-time business. They plan, each month, to share the profits in the ratio of the number of hours that each worked. During the first month, Rita worked 18 hours, Fred worked 30 hours, and Glen worked 12 hours. Express, in lowest terms, the ratio of the times they worked during the first month. Find the amount each should receive if the profit for this first month was $540.
Answer by checkley71(8403)

18+30+12=60 total hours worked.
Therefore Rita worked 18/60=3/10 the total hours.
Therefore Fred worked 30/60=5/10 or 1/2 the total hours.
Therefore Glen worked 12/60=2/10 or 1/5 the total hours.
The shares are 3/10, 5/10 & 2/10
Each share is worth 340/10=$34
Rita gets 3*34=$102
Fred gets 5*34=$170
Glen gets 2*34=$68
Adding up the amount of the 3 shares is the proof:
102+170+68=340
340=340